OpenStudy (anonymous):

Find the volume generated by revolving about the x-axis the area bounded by xy = 9, the x-axis and lines x = 3 and x=9

OpenStudy (amistre64):



OpenStudy (amistre64):

something similar to this right?

OpenStudy (anonymous):

Yess

OpenStudy (amistre64):

then we can integrate, or sum up the areas of circles, generated by the function from 3 to 9 i believe

OpenStudy (anonymous):

May I please see a full work out please?

OpenStudy (anonymous):

@ amistre64 Alright

OpenStudy (amistre64):

\[\sum_{3}^{4}\pi[f(x)]^2\ \Delta x\] \[\pi \int_{3}^{9}(\frac{9}{x})^2dx\]

OpenStudy (amistre64):